OpenStudy (anonymous):

sq rt 6m divided by sq rt 5? ;)

OpenStudy (fibonaccichick666):

is this \[\frac{\sqrt{6m}}{\sqrt{5}}\]correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (fibonaccichick666):

uhm, well, what do you want to do to it?

OpenStudy (anonymous):

\[\frac{\sqrt{30m} }{ 5 }\]

OpenStudy (anonymous):

hah you're doing PLATO too,huh?

OpenStudy (fibonaccichick666):

oh, rationalizing